OpenStudy (anonymous):

ok what is 14. Solve the area formula for a trapezoid, , for .

OpenStudy (anonymous):

it is A=1/2h(b1+b2)for b1

OpenStudy (anonymous):

A=0.5*h*(b1+b2) 2A/h=b1+b2 b1=2A/h-b2

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

Well, from your partially written out question, I assumed you want to solve the trapezoid formula for b1.

OpenStudy (anonymous):

no that is the question that my techer gave me

OpenStudy (anonymous):

So what is your question?

OpenStudy (anonymous):

14. Solve the area formula for a trapezoid, , for . (2 points) b1 = 2Ah – b2 1.b1= 2Ahb2 2.b1=2A/hb1 3.b1=1/2ahb1 4.b1= 2Ahb2 wech wo is rifgt

OpenStudy (anonymous):

i mean right

OpenStudy (anonymous):

None of them are right...

OpenStudy (anonymous):

yes i got it it is b1=AH-b2

OpenStudy (anonymous):

i mean it hase to be

OpenStudy (anonymous):

Why? Try going from that back to the area formula for a trapezoid.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i got b2=1/2hb1a

OpenStudy (anonymous):

Write out your steps? I swear you need a subtraction in there somewhere.

OpenStudy (anonymous):

me to

OpenStudy (anonymous):

b1= 1/2AH-b2

OpenStudy (anonymous):

i do not get any of it

OpenStudy (anonymous):

can you help

OpenStudy (sandra):

hey taking a look

OpenStudy (sandra):

so what was given was the first line? b1 = 2Ah – b2

OpenStudy (anonymous):

14. Solve the area formula for a trapezoid

OpenStudy (sandra):

ah ok so this is the formula of the trapazoid? t is A=1/2h(b1+b2)for b1

OpenStudy (sandra):

2A = 1/2h(b1+b2)

OpenStudy (sandra):

err sorry 2A/h = b1+ b2

OpenStudy (sandra):

b1 = 2A/h - b2

OpenStudy (anonymous):

but how

OpenStudy (anonymous):

that is what i do not get

OpenStudy (sandra):

ok so 1. A=1/2h(b1+b2) 2. A = h/2(b1 + b2) (another way to write 1/2h is h/2)

OpenStudy (sandra):

now we want to multiply both sides by 2

OpenStudy (sandra):

and we get 3. 2A = h(b1 + b2)

OpenStudy (sandra):

now we want to divide both sides by h, to get rid of the h - the whole point here is trying to get b1 by itself - so we continuously perform the same operation to both sides of the equation until it's alone - since if you do the same thing be it multiple, divide, or add to both sides, you still have a valid equation

OpenStudy (sandra):

so 2A/h = h(b1+b2)/h

OpenStudy (sandra):

and since we know h/h = 1, we can get rid of h on the right side

OpenStudy (sandra):

so 2A/h = b1+b2

OpenStudy (sandra):

so now, to get b1 by itself, we need to subtract b2 from both sides

OpenStudy (sandra):

so 2A/h - b2 = b1+b2-b2

OpenStudy (sandra):

that eliminates the b2 on the right side, and leaves b1 alone

OpenStudy (anonymous):

oh ok

OpenStudy (sandra):

sp 2A/h - b2 = b1, or b1 = 2A/h - b2

OpenStudy (anonymous):

ok i am starting to get it

OpenStudy (sandra):

but as far as the options go on your multiple choice - I don't see any that match

OpenStudy (anonymous):

ok i will contact my techer

OpenStudy (anonymous):

tomorrow

OpenStudy (sandra):

ok sorry =(

OpenStudy (sandra):

you could try inputting different values for b1 and b2

OpenStudy (sandra):

and h

OpenStudy (sandra):

and see which answer matches the from the choices and the original equation

OpenStudy (anonymous):

ok

OpenStudy (sandra):

but I would do it with our answer first

OpenStudy (sandra):

and see if they match up to the original area equation

OpenStudy (sandra):

if different values of b1,h, and b2 are always equal for our answer, it's almost certianly correct

OpenStudy (sandra):

and yeah, ask your teacher. worst case we're wrong and you have bonus points

OpenStudy (anonymous):

ok that would be cool lol

OpenStudy (anonymous):

thx for helping me